OpenStudy (anonymous):

rationalize the denominator and simplify 5+√5/8-√5

OpenStudy (anonymous):

\[\frac{5+\sqrt{5}}{8-\sqrt{5}}\] \[=\frac{5+\sqrt{5}}{8-\sqrt{5}}\times\frac{8+\sqrt{5}}{8+\sqrt{5}}=\frac{45+13\sqrt{5}}{59}\]